# Dataset Distillation with Probabilistic Latent Features

### v1

@@ Abstract
keywords:
Dataset Distillation , Stochastic Method , Image Classification
journal: Image and Vision Computing
[…]
[1]organization=IDEA lab, Friedrich Alexander Universität Erlangen-Nürnberg,addressline=Nürnberger Str. 74, city=Erlangen, postcode=91052, state=Bayern, country=Germany
[…]
[2]organization=Imperial College London,addressline=180 Queen’s Gate, city=London, postcode=SW72AZ, state=, country=UK
[…]
[3]organization=Technical University Munich,addressline=Trogerstrasse 18, city=Munich, postcode=81675, state=, country=Germany
@@ 3 Method
This section introduces the proposed Stochastic Latent Feature Distillation (SLFD) framework, with an overview provided in Fig.2. To offer an intuitive understanding of the approach, Fig.1 illustrates a simplified example involving two pixels within a probabilistic graphical model, highlighting the role of uncertainty and spatial correlation. The core of SLFD is a lightweight stochastic module (Fig.2($b$)), described in detail in Section3.2. This module learns a structured distribution over latent features, enabling the generation of diverse and spatially consistent synthetic samples during the distillation process. Section 3.3 and Fig. 2($c$) explain how we define a stochastic loss function that incorporates samples from the learned latent distribution. This loss guides the optimization process to produce synthetic data that more faithfully preserves the structure and variability of the original dataset. Finally, as shown in Fig. 2($d$), we assess the quality of the distilled dataset by training independent classifiers solely on the synthetic images.
$z_{1}$$z_{2}$$\eta_{1}$$y$$\eta_{2}$$x_{1}$$x_{2}$ $f_{\phi}(\boldsymbol{\eta})$
$z_{1}$$z_{2}$$\eta_{1}$$y$$\eta_{2}$$x_{1}$$x_{2}$ $f_{\phi}(\boldsymbol{\eta})$
Figure 1: Probabilistic graphical model for a two-pixel sample synthesis problem. Circular shapes indicate random variables, while diamond shapes represent deterministic variables. Variables that are shaded represent observed data, whereas those that are unshaded are unobserved. $\boldsymbol{x}$ are pixels of the synthesized image, $\boldsymbol{y}$ are the class labels, $\boldsymbol{\eta}$ are latent features, $f_{\phi}(\boldsymbol{\eta})$ indicates an implicit latent variable model, where $f_{\phi}$ can be a GAN’s generator. GAN discriminators are omitted as auxiliary objects that approximate an f-divergence. Note that our approach assumes $\boldsymbol{\eta}$ to be probabilistic and dependent in contrast to previous work [7]. We hypothesize that this allows to capture spatial correlations in latent space during the distillation phase.
@@ 3.2 Stochastic training
        Calculate gradient matching loss $\mathcal{L}=1-\frac{\nabla_{\theta}\mathcal{L}_{syn}^{c}\cdot\nabla_{\theta}\mathcal{L}_{real}^{c}}{\parallel\nabla_{\theta}\mathcal{L}_{syn}^{c}\parallel\parallel\nabla_{\theta}\mathcal{L}_{real}^{c}\parallel}$
@@
  Train five models only on the $ipc\times C$ synthetic images $S$.
@@ 3.3.1 Stochastic loss
$\mathcal{L}_{\text{syn}}^{c}=-\frac{1}{N_{w}\times M}\sum_{w=1}^{N_{w}}\sum_{m=1}^{M}\left(\log\sum_{m=1}^{M}\exp(y_{c}\cdot(q_{m}^{*})_{c})+\log M\right)$ (6)
Here, the label $y_{c}$ is repeated $M$ times to match the number of samples. As in prior work [10], we apply the log-sum-exp trick for numerical stability.
@@ 4.3.1 ImageNet
MTT [3] pixel $33.4_{\pm{1.5}}$ $34.0_{\pm{3.4}}$ $31.4_{\pm{3.4}}$ $27.7_{\pm{2.7}}$ $24.9_{\pm{1.8}}$ $24.1_{\pm{1.8}}$ $16.0_{\pm{1.2}}$ $25.5_{\pm{3.0}}$ $18.3_{\pm{2.3}}$ $18.7_{\pm{1.5}}$
GLaD [7] $39.9_{\pm{1.2}}$ $39.4_{\pm{1.3}}$ $34.9_{\pm{1.1}}$ $30.4_{\pm{1.5}}$ $29.0_{\pm{1.1}}$ $30.4_{\pm{1.5}}$ $17.1_{\pm{1.1}}$ $28.2_{\pm{1.1}}$ $21.1_{\pm{1.2}}$ $19.6_{\pm{1.2}}$
DC [1] pixel $38.7_{\pm{4.2}}$ $38.7_{\pm{1.0}}$ $33.3_{\pm{1.9}}$ $26.4_{\pm{1.1}}$ $27.4_{\pm{0.9}}$ $28.2_{\pm{1.4}}$ $17.4_{\pm{1.2}}$ $28.5_{\pm{1.4}}$ $20.4_{\pm{1.5}}$ $19.8_{\pm{0.9}}$
GLaD [7] $41.8_{\pm{1.7}}$ $42.1_{\pm{1.2}}$ $35.8_{\pm{1.4}}$ $28.0_{\pm{0.8}}$ $29.3_{\pm{1.3}}$ $31.0_{\pm{1.6}}$ $17.8_{\pm{1.1}}$ $29.1_{\pm{1.0}}$ $22.3_{\pm{1.6}}$ $21.2_{\pm{1.4}}$
SLFD $\mathbf{43.27_{\pm{1.58}}}$ $\mathbf{43.31_{\pm{1.13}}}$ $\mathbf{36.76_{\pm{1.10}}}$ $\mathbf{29.26_{\pm{1.46}}}$ $\mathbf{31.02_{\pm{1.40}}}$ $\mathbf{33.10_{\pm{1.18}}}$ $\mathbf{19.13_{\pm{1.05}}}$ $\mathbf{30.22_{\pm{1.22}}}$ $\mathbf{23.62_{\pm{1.26}}}$ $\mathbf{21.65_{\pm{1.30}}}$
GLaD-Conv $44.1_{\pm{2.4}}$ $49.2_{\pm{1.1}}$ $42.0_{\pm{0.6}}$ $35.6_{\pm{0.9}}$ $\mathbf{35.8_{\pm{0.9}}}$ $35.4_{\pm{1.2}}$ $22.3_{\pm{1.1}}$ $33.8_{\pm{0.9}}$ $20.7_{\pm{1.1}}$ $22.6_{\pm{0.8}}$
SLFD-Conv $\mathbf{46.00_{\pm{1.39}}}$ $\mathbf{49.96_{\pm{0.97}}}$ $\mathbf{42.60_{\pm{1.00}}}$ $\mathbf{36.40_{\pm{2.26}}}$ $35.44_{\pm{0.60}}$ $\mathbf{36.72_{\pm{1.31}}}$ $\mathbf{22.68_{\pm{0.92}}}$ $\mathbf{33.88_{\pm{1.01}}}$ $\mathbf{21.96_{\pm{0.45}}}$ $\mathbf{24.12_{\pm{1.03}}}$
DM [4] pixel $27.2_{\pm{1.2}}$ $24.4_{\pm{1.1}}$ $23.0_{\pm{1.4}}$ $18.4_{\pm{1.7}}$ $17.7_{\pm{0.9}}$ $20.6_{\pm{0.7}}$ $14.5_{\pm{0.9}}$ $17.8_{\pm{0.8}}$ $14.5_{\pm{1.1}}$ $14.0_{\pm{1.1}}$
GLaD [7] $31.6_{\pm{1.4}}$ $31.3_{\pm{3.9}}$ $26.9_{\pm{1.2}}$ $21.5_{\pm{1.0}}$ $20.4_{\pm{0.8}}$ $21.9_{\pm{1.1}}$ $15.2_{\pm{0.9}}$ $18.2_{\pm{1.0}}$ $20.4_{\pm{1.6}}$ $16.1_{\pm{0.7}}$
SLFD $\mathbf{39.18_{\pm{1.57}}}$ $\mathbf{38.30_{\pm{1.08}}}$ $\mathbf{32.02_{\pm{2.30}}}$ $\mathbf{25.21_{\pm{2.10}}}$ $\mathbf{23.52_{\pm{1.73}}}$ $\mathbf{28.67_{\pm{1.00}}}$ $\mathbf{17.53_{\pm{1.03}}}$ $\mathbf{24.70_{\pm{2.55}}}$ $\mathbf{23.25_{\pm{1.54}}}$ $\mathbf{18.37_{\pm{0.71}}}$
GLaD-Conv $41.0_{\pm{1.5}}$ $42.9_{\pm{1.9}}$ $39.4_{\pm{0.7}}$ $33.2_{\pm{1.4}}$ $30.3_{\pm{1.3}}$ $32.2_{\pm{1.7}}$ $21.2_{\pm{1.5}}$ $27.6_{\pm{1.9}}$ $21.8_{\pm{1.8}}$ $22.3_{\pm{1.6}}$
SLFD-Conv $\mathbf{58.72_{\pm{1.50}}}$ $\mathbf{54.04_{\pm{1.13}}}$ $\mathbf{46.56_{\pm{1.66}}}$ $\mathbf{41.88_{\pm{1.09}}}$ $\mathbf{37.68_{\pm{0.70}}}$ $\mathbf{43.72_{\pm{0.92}}}$ $\mathbf{25.44_{\pm{1.57}}}$ $\mathbf{38.08_{\pm{0.92}}}$ $\mathbf{35.44_{\pm{0.72}}}$ $\mathbf{27.56_{\pm{0.77}}}$
@@
SLFD $\mathbf{44.01_{\pm{1.68}}}$ $\mathbf{43.15_{\pm{1.64}}}$ $\mathbf{37.28_{\pm{1.38}}}$ $\mathbf{29.78_{\pm{1.07}}}$ $\mathbf{31.57_{\pm{0.99}}}$
@@
average AlexNet ResNet18 VGG11 ViT
SLFD (50) $14.28_{\pm{0.91}}$ $12.60_{\pm{1.41}}$ $11.02_{\pm{0.91}}$ $17.25_{\pm{0.64}}$ $16.23_{\pm{0.70}}$
SLFD (100) $7.47_{\pm{0.37}}$ $6.64_{\pm{0.36}}$ $4.18_{\pm{0.33}}$ $9.24_{\pm{0.42}}$ $9.81_{\pm{0.35}}$
Table 3: Cross architecture results with more classes (in brackets) with DC.
To further evaluate the scalability and generalization of our method, we extend our experiments beyond individual subsets. Specifically, we combine $5$ subsets (ImNet-A through ImNet-E), each with a resolution of $128\times 128$, to form a new dataset comprising $50$ classes. Additionally, we merge all $10$ subsets to create a $100$-class configuration. The cross-architecture performance of our method on these larger and more diverse datasets is reported in Table 3, where we evaluate using four classifiers that were not involved in the distillation process.
@@ 4.3.2 CIFAR-10
pixel $26.0_{\pm{0.4}}$ $25.9_{\pm{0.2}}$ $27.3_{\pm{0.5}}$ $28.0_{\pm{0.5}}$ $22.9_{\pm{0.3}}$
GLaD $G_{r}$ [7] $26.6_{\pm{0.6}}$ $\mathbf{30.1_{\pm{0.5}}}$ $27.3_{\pm{0.2}}$ $28.0_{\pm{0.9}}$ $21.2_{\pm{0.6}}$
GLaD $G_{t}$ [7] $26.3_{\pm{0.5}}$ $26.0_{\pm{0.7}}$ $27.6_{\pm{0.6}}$ $28.2_{\pm{0.6}}$ $23.4_{\pm{0.2}}$
SLFD $G_{t}$ $\mathbf{28.1_{\pm{0.55}}}$ $26.79_{\pm{0.55}}$ $\mathbf{28.45_{\pm{0.94}}}$ $\mathbf{28.73_{\pm{0.38}}}$ $\mathbf{28.45_{\pm{0.33}}}$
@@ 4.4 Ablation study
GAN GLaD $37.4_{\pm{5.5}}$ $41.5_{\pm{1.2}}$ $35.7_{\pm{4.0}}$ $27.9_{\pm{1.0}}$ $29.3_{\pm{1.2}}$
SLFD $\mathbf{44.01_{\pm{1.68}}}$ $\mathbf{43.15_{\pm{1.64}}}$ $\mathbf{37.28_{\pm{1.38}}}$ $\mathbf{29.78_{\pm{1.07}}}$ $\mathbf{31.57_{\pm{0.99}}}$
UViT GLaD $37.70_{\pm{1.53}}$ $38.96_{\pm{1.68}}$ $\mathbf{33.11_{\pm{1.65}}}$ $26.38_{\pm{1.24}}$ $27.34_{\pm{1.11}}$
[…]
@@
128 GLaD(DC) [7] $41.8_{\pm{1.7}}$ $42.1_{\pm{1.2}}$ $35.8_{\pm{1.4}}$ $28.0_{\pm{0.8}}$ $29.3_{\pm{1.3}}$
SLFDrank=10 $\mathbf{43.27_{\pm{1.58}}}$ $\mathbf{43.31_{\pm{1.13}}}$ $36.76_{\pm{1.10}}$ $29.26_{\pm{1.46}}$ $31.02_{\pm{1.40}}$
SLFDmulti $43.21_{\pm{1.55}}$ $43.20_{\pm{1.13}}$ $\mathbf{36.81_{\pm{1.57}}}$ $\mathbf{29.50_{\pm{1.09}}}$ $\mathbf{31.25_{\pm{1.31}}}$
256 GLaD(DC) [7] $37.4_{\pm{5.5}}$ $41.5_{\pm{1.2}}$ $35.7_{\pm{4.0}}$ $27.9_{\pm{1.0}}$ $29.3_{\pm{1.2}}$
SLFDrank=10 $44.01_{\pm{1.68}}$ $43.15_{\pm{1.64}}$ $37.28_{\pm{1.38}}$ $29.78_{\pm{1.07}}$ $31.57_{\pm{0.99}}$
SLFDmulti $\mathbf{44.16_{\pm{1.21}}}$ $\mathbf{43.39_{\pm{1.44}}}$ $\mathbf{37.40_{\pm{1.18}}}$ $\mathbf{30.41_{\pm{1.16}}}$ $\mathbf{32.05_{\pm{1.36}}}$
@@ Acknowledgments
Generative AI tools were used solely to assist with word choice and to improve the language and readability of the manuscript. No content was generated or interpreted by AI beyond these purposes.

### v2

@@ 3 Method
This section introduces the proposed Stochastic Latent Feature Distillation (SLFD) framework, with an overview provided in Fig.1. The core of SLFD is a lightweight stochastic module (Fig.1($b$)), described in detail in Section3.2. This module learns a structured distribution over latent features, enabling the generation of diverse and spatially consistent synthetic samples during the distillation process. Section 3.3 and Fig. 1($c$) explain how we define a stochastic loss function that incorporates samples from the learned latent distribution. This loss guides the optimization process to produce synthetic data that more faithfully preserves the structure and variability of the original dataset. Finally, as shown in Fig. 1($d$), we assess the quality of the distilled dataset by training independent classifiers solely on the synthetic images.
@@ 3.1 Preliminaries
        Gradient matching $\mathcal{L}=1-\frac{\nabla_{\theta}\mathcal{L}_{syn}^{c}\cdot\nabla_{\theta}\mathcal{L}_{real}^{c}}{\parallel\nabla_{\theta}\mathcal{L}_{syn}^{c}\parallel\parallel\nabla_{\theta}\mathcal{L}_{real}^{c}\parallel}$
@@
  Train classifiers only on the synthetic images $S$.
@@ 3.3.1 Stochastic loss
$\mathcal{L}_{\text{syn}}^{c}=-\frac{1}{L}\sum_{w=1}^{N_{w}}\sum_{m=1}^{M}\left(\log\sum_{m=1}^{M}\exp(y_{c}\cdot(q_{m}^{*})_{c})+\log M\right)$ (6)
Here, the label $y_{c}$ is repeated $M$ times to match the number of samples, $L=N_{w}\times M$. As in prior work [23], we apply the log-sum-exp trick for numerical stability.
@@ 3.3.2 Gradient matching loss
MTT [3] pixel 33.4±1.5 34.0±3.4 31.4±3.4 27.7±2.7 24.9±1.8 24.1±1.8 16.0±1.2 25.5±3.0 18.3±2.3 18.7±1.5
GLaD [4] 39.9±1.2 39.4±1.3 34.9±1.1 30.4±1.5 29.0±1.1 30.4±1.5 17.1±1.1 28.2±1.1 21.1±1.2 19.6±1.2
DC [39] pixel 38.7±4.2 38.7±1.0 33.3±1.9 26.4±1.1 27.4±0.9 28.2±1.4 17.4±1.2 28.5±1.4 20.4±1.5 19.8±0.9
GLaD [4] 41.8±1.7 42.1±1.2 35.8±1.4 28.0±0.8 29.3±1.3 31.0±1.6 17.8±1.1 29.1±1.0 22.3±1.6 21.2±1.4
SLFD 43.27±1.6 43.31±1.1 36.76±1.1 29.26±1.5 31.02±1.4 33.10±1.2 19.13±1.1 30.22±1.2 23.62±1.3 21.65±1.3
GLaD-Conv 44.1±2.4 49.2±1.1 42.0±0.6 35.6±0.9 35.8±0.9 35.4±1.2 22.3±1.1 33.8±0.9 20.7±1.1 22.6±0.8
SLFD-Conv 46.00±1.4 49.96±1.0 42.60±1.0 36.40±2.3 35.44±0.60 36.72±1.3 22.68±0.9 33.88±1.0 21.96±0.5 24.12±1.0
DM [38] pixel 27.2±1.2 24.4±1.1 23.0±1.4 18.4±1.7 17.7±0.9 20.6±0.7 14.5±0.9 17.8±0.8 14.5±1.1 14.0±1.1
GLaD [4] 31.6±1.4 31.3±3.9 26.9±1.2 21.5±1.0 20.4±0.8 21.9±1.1 15.2±0.9 18.2±1.0 20.4±1.6 16.1±0.7
SLFD 39.18±1.6 38.30±1.1 32.02±2.3 25.21±2.1 23.52±1.7 28.67±1.0 17.53±1.0 24.70±2.6 23.25±1.5 18.37±0.7
GLaD-Conv 41.0±1.5 42.9±1.9 39.4±0.7 33.2±1.4 30.3±1.3 32.2±1.7 21.2±1.5 27.6±1.9 21.8±1.8 22.3±1.6
SLFD-Conv 58.72±1.5 54.04±1.1 46.56±1.7 41.88±1.1 37.68±0.7 43.72±0.9 25.44±1.6 38.08±0.9 35.44±0.7 27.56±0.8
@@ 4.3.1 ImageNet
SLFD 44.01±1.7 43.15±1.6 37.28±1.4 29.78±1.1 31.57±1.0
@@ 4.3.2 CIFAR-10
pixel 26.0±0.4 25.9±0.2 27.3±0.5 28.0±0.5 22.9±0.3
GLaD $G_{r}$ [4] 26.6±0.6 30.1±0.5 27.3±0.2 28.0±0.9 21.2±0.6
GLaD $G_{t}$ [4] 26.3±0.5 26.0±0.7 27.6±0.6 28.2±0.6 23.4±0.2
SLFD $G_{t}$ 28.1±0.6 26.8±0.6 28.5±0.9 28.7±0.4 28.5±0.3
@@ 4.4 Ablation study
GAN GLaD 37.4±5.5 41.5±1.2 35.7±4.0 27.9±1.0 29.3±1.2
SLFD 44.01±1.7 43.15±1.6 37.28±1.4 29.78±1.1 31.57±1.0
UViT GLaD 37.70±1.5 38.96±1.7 33.11±1.7 26.38±1.2 27.34±1.1
SLFD 37.92±2.1 38.97±1.8 32.67±1.5 26.78±1.3 28.21±1.6
@@
128 GLaD(DC) [4] 41.8±1.7 42.1±1.2 35.8±1.4 28.0±0.8 29.3±1.3
SLFDrank=10 43.27±1.6 43.31±1.1 36.76±1.1 29.26±1.5 31.02±1.4
SLFDmulti 43.21±1.6 43.20±1.1 36.81±1.6 29.50±1.1 31.25±1.3
256 GLaD(DC) [4] 37.4±5.5 41.5±1.2 35.7±4.0 27.9±1.0 29.3±1.2
SLFDrank=10 44.01±1.7 43.15±1.6 37.28±1.4 29.78±1.1 31.57±1.0
SLFDmulti 44.16±1.2 43.39±1.4 37.40±1.2 30.41±1.2 32.05±1.4
@@ Acknowledgments
Supplementary Material
6 Rationale
Having the supplementary compiled together with the main paper means that:
  • The supplementary can back-reference sections of the main paper, for example, we can refer to Sec. 1;
  • The main paper can forward reference sub-sections within the supplementary explicitly (e.g. referring to a particular experiment);
  • When submitted to arXiv, the supplementary will already included at the end of the paper.
To split the supplementary pages from the main paper, you can use Preview (on macOS), Adobe Acrobat (on all OSs), as well as command line tools.